### v1

# UBGAN: Enhancing Coded Speech with Blind and Guided Bandwidth Extension

Abstract
In practical application of speech codecs, a multitude of factors such as the quality of the radio connection, limiting hardware or required user experience necessitate trade-offs between achievable perceptual quality, engendered bitrate and computational complexity. Most conventional and neural speech codecs operate on wideband (WB) speech signals to achieve this compromise. To further enhance the perceptual quality of coded speech, bandwidth extension (BWE) of the transmitted speech is an attractive and popular technique in conventional speech coding. In contrast, neural speech codecs are typically trained end-to-end to a specific set of requirements and are often not easily adaptable. In particular, they are typically trained to operate at a single fixed sampling rate. With the Universal Bandwidth Extension Generative Adversarial Network (UBGAN), we propose a modular and lightweight GAN-based solution that increases the operational flexibility of a wide range of conventional and neural codecs. Our model operates in the subband domain and extends the bandwidth of WB signals from 8 kHz to 16 kHz, resulting in super-wideband (SWB) signals. We further introduce two variants, guided-UBGAN and blind-UBGAN, where the guided version transmits quantized learned representation as a side information at a very low bitrate additional to the bitrate of the codec, while blind-BWE operates without such side-information. Our subjective assessments demonstrate the advantage of UBGAN applied to WB codecs and highlight the generalization capacity of our proposed method across multiple codecs and bitrates.
1 Introduction
Speech codecs for real-time communications have evolved significantly over the years. With the increasing demand for higher speech quality, the limitations of traditional codecs, such as Adaptive Multi-Rate narrowband (AMR-NB[1], became apparent due to their limited bandwidth. This led to the adoption of more advanced codecs, including AMR-WB [1], a WB codec with sampling rate of 16 kHz and Enhanced Voice Services (EVS[2], which supports multiple bandwidths, including a SWB mode with a sampling rate of 32 kHz. A key aspect of this evolution has been the adoption of BWE, a technique that generates high-frequency content of speech signals from the received coded low-frequency components and possibly from other transmitted parameters. This approach is widely used alongside core codecs to enhance the quality of the decoded signals. BWE enables codecs to allocate bits more effectively and produces more intelligible speech through the reconstruction of high-frequency content.
BWE has become a pivotal part of modern classical speech codecs, which is usually employed as blind-BWE (BBWE) for the lowest bitrates or as guided-BWE (GBWE) utilizing side-information when the available bit-budget is higher. AMR-WB features a Code-Excited Linear Prediction (CELP) core that operates in the 0-6.4 kHz band and applies BBWE to extend the bandwidth up to 7 kHz [3] by shaping a white noise signal with a interpolated Linear Prediction (LP) synthesis filter. More advanced codecs like EVS offer a CELP core at two different internal sampling rates: 12.8 kHz and 16 kHz depending on the bitrates. It employs a more sophisticated technique, known as time domain bandwidth extension (TBE[4] to obtain either WB or SWB signals. In TBE, the high frequency fine structure (excitation) is generated by applying a non-linear operation on the low-band excitation before being spectrally shaped with an LP synthesis filter that is sent as side information along with energy gains.
While classical codecs are still widely used, neural speech codecs have been gaining momentum in recent years as they are capable of generating high quality speech at very low bitrates. The earliest approaches [5][6][7] typically focused on decoding speech using signal features such as Mel Frequency Cepstral Coefficients (MFCCs), Mel-spectrograms or other acoustic parameters, and were limited to generating WB signals due to the low resolution of these representations at higher frequency bands. The rise of end-to-end trained neural codecs improved achievable quality and enabled the support of a wide range of bandwidth such as WB[8][9], semi-SWB (24 kHz sampling rate)[10][11][12] and even Fullband (FB) (48 kHz sampling rate)[13]. However, most of these end-to-end systems rely on time-domain auto-encoder architectures that use learned upsampling and downsampling layers, making the computational complexity of the model proportional to the output sampling rate. More recently, short-time Fourier transform (STFT)-based codecs[14][15][16][17] provide a more reliable way to work with different bandwidths without exploding the complexity of the model but still suffer difficulties at capturing finer details of high-frequency content of the signal.
Although neural codecs trained end-to-end offer promising solutions, the importance of BWE cannot be overstated: It offers a modularity and scalability in terms of computational complexity, bitrate and signal quality that is attractive for many applications. A very similar approach to classical methods was proposed in [18] that used a neural codec for the WB signal component and a classical transform-domain BWE method to code higher bands. Neural networks have been widely explored for BWE as well and have also commonly been referred to as audio super-resolution methods. These models either operate in the time domain [19][20][21] or in the frequency domain [22][23][24] and are designed to either regenerate the entire signal or to only generate the missing frequency bands. These models are capable of performing high-quality BWE but come at the cost of higher computational complexity and often inject additional algorithmic delay. Even the streaming capable models [25][26] are highly complex when considered for hand-held devices. Also, most research in this area is limited to expanding the bandwidth of non-coded speech signals.
Coded speech signals undergo severe degradation, especially at low bitrates. For best signal quality, these degradations need to be considered when generating high-frequency content from the low-frequency content. Some approaches [27][28][29] have been developed particularly for conventional speech codecs like AMR-NB and Opus[30]. However, enhancement of multiple speech codecs, including neural ones, with a single generalized BWE model remains largely unexplored. Furthermore, to the best of our knowledge, guided BWE using neural networks has not yet been investigated. In this paper, we present UBGAN, a GAN-based BWE model that is trained to work with multiple WB codecs to generate SWB speech signals. Our contribution in this paper can be summarized as below:
  • We propose a causal, lightweight BWE model capable of extending WB coded speech to SWB for both conventional and neural speech codecs.
  • We propose robust high band synthesis in the Pseudo-Quadrature Mirror Filters (PQMF) subband domain, i.e., the proposed model is capable of generating high-frequency content even for a varying level of degradation in the low-frequency content.
  • We propose blind-UBGAN that is a BWE model conditioned on the mel-spectrogram calculated from the lower frequency bands of the coded speech signal.
  • We also propose guided-UBGAN, which relies on additionally transmitted quantized learned features from the upper PQMF subbands for better steering of the BWE system. The quantized features require an additional bitrate of 0.2 kbps.
2 Proposed Model
A schematic diagram of UBGAN is shown in Fig. 1. The proposed model can be structured into three primary blocks described in the following sections: the conditioning block, the synthesis block, and an overlap-compensation block.
2.1 Conditioning Block
The conditioning block calculates features for conditioning the synthesis block. We use an 80-band Mel-spectrogram as conditioning for the blind-UBGAN model. It is calculated from the coded signal on a frame size of 20 ms with 5 ms of past context and 5 ms of look-ahead to effectively capture temporal dependencies.
In contrast, the guided-UBGAN model employs a neural side-info encoder block that is trained in an end-to-end manner alongside the synthesis block. This integrated approach enhances representation learning and synthesis quality. The side-info encoder consists of four key components: an eight-band analysis PQMF, a rolling window mechanism, a Dual-Path Convolutional Recurrent Network (DPCRNN) block [9] and a scalar quantization (SQ) module [31]. The input signal at sampling rate of 32 kHz, undergoes PQMF analysis, resulting in eight subbands, each with 2 kHz bandwidth. To focus on higher frequency components, we discard the lowest four bands containing low-frequency signal components. The rolling window reshapes the data such that each frame consists of concatenated 20 ms segments from each of the higher PQMF bands, accompanied by a 5 ms history and a 5 ms look-ahead. Each frame goes through a DPCRNN block that consists of a Gated Recurrent Unit (GRU) sandwiched between two $1\times 1$ convolution layers. The final output of the DPCRNN is an 80-dimensional latent vector per frame which is then projected into a single dimension using a fully-connected layer before being quantized using 4 bits. Hence, GBWE requires 0.2 kbps of additional bitrate.
Figure 1: High-level schematics of UBGAN. Dashed line represents residual connection from downsample blocks to upsample blocks.
Figure 2: Description of the downsample and upsample block of UBGAN
2.2 Synthesis Block
The synthesis block is the pivotal part of the system. It works in the time domain and takes the coded signal along with the conditioning information as input. Initially, the WB coded speech is decomposed into four subbands using PQMF analysis. The synthesis block then predicts additional four subbands, extending the bandwidth through an eight-band PQMF synthesis.
The architecture of the model is heavily inspired by PostGAN [32], which comprises several key components, including PQMF analysis and pre-conv, post-conv, downsample as well as upsample blocks. The pre and post-conv blocks contain a single convolutional layer with kernel size of 7 that maps the four input channels to eight output channels and vice-versa. The downsample and upsample blocks create a UNet-like structure with residual connections as shown in Fig. 2.
In the downsample block, the input goes through channel normalization, 1-D convolution, gated activation and an interpolation layer sandwiched between two 1D convolutions [32]. We use six blocks with a varying number of channels [8, 16, 32, 48, 48, 64]. The downsampling of the input signal is done in each block via interpolation by a factor of [1, 2, 2, 2, 2.5, 2], respectively. The upsampling of the conditioning feature is done with scaling factors [80, 80, 40, 20, 10, 4] and the result is passed through a convolution layer. We obtain two 64-dimensional latent vectors for each 20 ms frame after the last downsample block. The output before the gated activation is merged with the upsampled conditioning signal to obtain the Temporal Adaptive DE-normalization (TADE) parameters $\gamma$ and $\beta$ for the residual connection. Each upsample block contains a TADE-residual block similar to [6]. The TADE parameters from the corresponding downsample block is used for performing an affine transformation of the normalized input. The operations normalization, linear transformation and gated activation are each applied twice before the output is upsampled. The sequences of channel numbers and upsampling scale factors are the reverse of the downsample block parameters. We use causal depthwise-separable convolutions (DSConvs) in all the layers and a kernel size of 7. The use of DSConvs, along with a varying number of channels in the upsample and downsample blocks helps to achieve a model with low computational complexity and fewer parameters.
2.3 Overlap Compensation Block
In our PQMF subband processing framework, the input coded speech is initially decomposed into four subbands, each with a bandwidth of 2 kHz. Subsequently, the output speech is synthesized using eight PQMF subbands, with the four highest subbands being generated by the network. A simple concatenation of input and generated subbands produces audible artefacts due to overlap of the subbands and the inherent aliasing of the PQMF decomposition. Specifically, the transition between the fourth and fifth subband requires extra effort to ensure that the aliasing between these two subbands cancels out effectively. To address this issue, we propose an overlap compensation block, which comprises two convolutional layers with tanh activation function. This block takes the highest subband of the coded speech and the lowest subband of the model-generated output, and outputs those subband signals with well-compensated overlap. This ensures a seamless transition during synthesis and eliminates potential artifacts.
2.4 Discriminator
For adversarial training, we use an ensemble of four STFT discriminators proposed in [11]. The discriminators of the ensemble are working on multiple window lengths (2048, 1024, 512 and 256) with 75% overlap. We also tried the multi-scale discriminator of [15][9] but found that such time-domain discriminators do not perform as well as their frequency-domain counterpart for SWB signals.
3 Experimental Setup
3.1 Training and Losses
We trained the models in a two-step procedure similar to [9] [15] [32]. First, a pre-training is done for 50 epochs, then an adversarial training phase is started. Let $x$ and $\hat{x}$ denote the target and generated SWB signals and $|X|$ and $|\hat{X}|$ their corresponding STFT magnitudes. The pre-training loss consists of a spectral convergence loss
$\mathcal{L}_{\text{sc}}=\frac{\left\|\,|X|-|\hat{X}|\,\right\|_{F}}{\left\|\,|X|\,\right\|_{F}}$ (1)
and a log-magnitude loss
$\mathcal{L}_{\text{mag}}=\frac{1}{N}\left\|\,\log(|X|)-\log(|\hat{X}|)\,\right\|_{1},$ (2)
where .$\left\|\,.\,\right\|_{F}$, .$\left\|\,.\,\right\|_{1}$ and $N$ represents the Frobenius and the L1 norms and the number of time-frequency bins, respectively. As adversarial loss for training the generator we use the Least-Squares GAN (LSGAN) loss [33]
$\mathcal{L}_{\text{adv}}=\mathbb{E}_{p(x)}\left[\sum_{k=1}^{4}\left(D_{k}(G(x))-1\right)^{2}\right],$ (3)
where $D_{k}$ represents the $k^{th}$ discriminator for generated signal $G(x)$, and $p(x)$ denote the probability density of the training data $x$. Alongside the LSGAN loss, we use a feature loss that calculates the L1-loss between features from the generated signal and the ground truth signal at the output of each discriminator layer. Finally, the feature loss is scaled by a factor of 10 while all other losses are added unscaled.
For generator training, we use the ADAMW optimizer with learning rate $\text{lr}_{\text{g}}=5e\text{-}4$ and exponential decay by a factor of 0.99 every five epochs. The discriminator is trained with the ADAM optimizer with learning rate $\text{lr}_{\text{d}}=2e\text{-}4$. We use a batch size of 64 and the straight-through estimator is used at the quantizer bottleneck for training the side-info encoder of guided-UBGAN.
3.2 Dataset
The model was trained with 562 hours of the read speech dataset [34] and 44 hours of the VCTK dataset [35]. Both datasets contain FB signals which are resampled to 16 kHz for input and 32 kHz for SWB output signals. The model is trained with coded speech from two conventional codecs and two neural codecs. We encode and decode each 16 kHz signal using the following codecs with equal probability: EVS 5.9 kbps [2], Opus 6 kbps [30], LyraV2 3.2 kbps [10], STFTNC 1.5 kbps [15]. For both subjective and objective evaluation, we select a diverse set of 24 items created from various sources including ITU-T Standard P.501 [36], ETSI TS 103 281 Annex E [37] and NTT [38].
3.3 Evaluation
For evaluation, we carry out both subjective and objective evaluations. For objective assessment, we use the 2f-Model [39] and the VISQOLAudio V3 metric [40] as they both support evaluation over FB speech signals. For subjective assessment, we conduct a P.808 Degradation Category Ratings (DCR) listening test [41] using the Amazon Mechanical Turk service.
3.4 Baselines
In our evaluation, we employed a diverse set of codecs at various bitrates, which we classify in the following categories:
  • Seen WB codecs: LyraV2 (3.2 kbps) [10], STFTNC (1.5 kbps) [15], EVS (7.2 kbps) [2], Opus (9 kbps) [30]. Note that the bitrates used for EVS and Opus are different than in training.
  • Unseen WB codecs: Funcodec (3 kbps) [14], LPCNet (1.6 kbps) [7], AMR-WB (6.6 kbps) [3].
  • Semi-SWB/SWB codecs: DAC IBM (1.5 kbps) [42], Encodec (3 kbps) [11], EVS-SWB (9.6 kbps) [4]. The outputs of these models were resampled to 16 kHz before BWE.
  • BWE-model: We use the STFT-based AP-BWE [24] model as baseline. It performs FB extension of the WB speech up to a bandwidth of 48 kHz.
4 Results and Discussion
[…]
WB Blind Guided AP-BWE WB Blind Guided AP-BWE
AMR-WB (6.6 kbps) $19.31$ $31.55$ 34.64 $22.87$ $2.79$ $3.33$ 3.54 $2.64$
EVS (7.2 kbps) $20.85$ $31.50$ 36.26 $25.30$ $2.86$ $3.28$ 3.59 $2.88$
Opus (9 kbps) $28.64$ $39.48$ 41.40 $35.21$ $2.95$ $3.56$ 3.64 $3.02$
FunCodec (3 kbps) $24.28$ $35.06$ 38.52 $29.78$ $2.88$ $3.42$ 3.55 $2.94$
LPCNet (1.6 kbps) $23.20$ $32.29$ 33.49 $27.67$ $2.72$ $3.05$ 3.36 $2.87$
LyraV2 (3.2 kbps) $25.42$ $36.21$ 37.51 $31.59$ $2.86$ $3.48$ 3.53 $3.07$
STFTNC (1.5 kbps) $23.06$ $31.63$ 34.38 $27.54$ $2.84$ $3.36$ 3.50 $2.97$
Table 1: Average VISQOL-Audio and 2f-Model scores of our proposed method and the baseline. Blind and guided denote the resulting SWB signals from blind- and guided-UBGAN, respectively. AP-BWE is the baseline BWE method for obtaining FB speech. For all conditions, the FB signals are used as reference.
In Table 1, we present the objective scores obtained for various codecs with our proposed and baseline BWE methods. Both objective measures are reference-based, thus for uniformity, for each measure, the FB signals are used as reference irrespective of the bandwidth of the signals under test. The 2f-Model as well as the VISQOL scores show the quality improvement obtained through BWE over WB speech. For blind- and guided-UBGAN, both metrics show the superiority of the guided model which is able to reconstruct higher frequencies better. Despite higher bandwidth, the AP-BWE method scores lower than both of our proposed methods. Upon closer inspection, we found that the baseline method AP-BWE that was trained on clean speech injects the coding artefacts from the lower bands to higher bands, which degrades the quality of the resulting FB signal.
For a better assessment of our proposed method, we conducted two P.808 DCR subjective listening tests. DCR tests include the presentation of the reference signal to the listeners before the degraded items are presented. In the first test, we evaluated the benefit of our proposed BWE method over WB codecs. The results shown in Fig. 3 clearly indicate the improvement in quality brought by wider bandwidth. The guided-UBGAN performs slightly better than the blind variant for all codecs except for Funcodec and OPUS. We investigated the individual scores of the items for these codecs and found that, most items in the guided case are rated almost on par with or above the blind-UBGAN except for a few stimuli where degradation of the coded signals seems to be substantial. Upon casual listening, we established that the core-codec quality has a major impact on the perceived quality of the extended bandwidth. Here, the guided BWE version can lead to a mismatch between the energy levels of the extended bandwidth components and the coded core bandwidth components. This is because this model is conditioned by the encoded high frequency side-information derived from the clean signal, which might be of different energy than the coded signals.
Figure 3: P.808 DCR scores with 27 listeners for WB codecs with blind-UBGAN and guided-UBGAN.
Figure 4: P.808 DCR scores with 21 listeners for semi-SWB/SWB codecs with guided-UBGAN and AP-BWE baseline
The results shown in Fig. 4 compare semi-SWB/SWB codecs, our proposed guided-UBGAN, and the output of the pre-trained AP-BWE model capable of full-band extension of the signal. The core-coded signals are resampled to 16 kHz before performing BWE and the used reference signal has a sampling rate of 48 kHz. For the neural codecs with 12 kHz bandwidth, our proposed approach that reconstructs the lost bandwidth due to resampling, performs similar to the semi-SWB baseline codecs themselves, thus showing the effectiveness of our modular approach compared to the end-to-end approaches. However, for the conventional SWB codec EVS, the proposed approach falls short. This may be attributed to the fact that the parametric-BWE in EVS uses a higher bitrate of 0.95 kbps to perform guided BWE and is highly tuned to the core-coder and the given operating point. Also, our proposed model was trained on lower bitrates and lower signal qualities of core-codecs and, thus, has limitations for codecs achieving higher signal quality. For almost all codecs, the computationally complex AP-BWE method improves signal quality relative to the codec output and the proposed solution, but the perceived quality difference is not substantial as can be seen in Fig. 4. From our results, we also observe that the FB extension of speech signal might not lead to substantial quality improvements and semi-SWB/SWB or WB codecs associated with (semi) SWB-BWE can be seen as efficient design choices, especially in the paradigm of neural codecs, where model complexity increases with the core-bandwidth.
Models Complexity (GFLOPS) Parameter Count
AP-BWE (16 kHz to 48 kHz) 17.87 29.76M
Blind-UBGAN 0.214 250K
Guided-UBGAN 0.246 364K
Table 2: Complexity and parameter count of the proposed model and baseline
The computational complexity and the parameter count of the compared models are shown in Table 2. Our proposed blind- and guided-UBGAN models are significantly less complex than the AP-BWE and also have fewer parameters. The resulting computational efficiency enables the model to run on devices with low computational power in real time and also facilitates their integration with other neural codecs while maintaining minimal complexity overhead. Furthermore, the proposed models are causal and operate on a frame size of 20 ms and a look-ahead of 5 ms, making them suitable for integration into a variety of codecs without introducing significant algorithmic delay.
5 Conclusion
In this paper, we introduced a universal low-complex, low-delay and causal BWE method for WB speech codecs. We demonstrate its effectiveness across a wide range of speech codecs. The proposed model works in the PQMF domain and synthesizes the higher subbands of the signal. Furthermore, we introduced a guided version of it that learns and quantize high-frequency features at a low bitrate of 0.2 kbps from the original speech. Our findings indicate that the guided-UBGAN model offers significantly greater quality improvements than the blind-UBGAN for most codecs. Additionally, our analysis highlights the importance of bandwidth in perceived speech quality. It also emphasizes that for low-bitrate codecs, the quality of the speech signal tends to saturates at semi-SWB or SWB and the advantage of performing FB extension might be limited. In future work, we intend to extend the model to general audio applications and investigate ways to improve the performance of guided methods.
6 Acknowledgements
We would like to thank Markus Multrus and Anika Treffehn from Farunhofer IIS for their support. Parts of this work have been funded by the Free State of Bavaria in the DSgenAI project. We thankfully acknowledge the scientific support and HPC resources provided by the Erlangen National High Performance Computing Center (NHR@FAU).